OpenStudy (anonymous):

(-2,2) (5,3) find the slope

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 2}})\quad &({\color{red}{ -5}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (akashdeepdeb):

Slope of a given line if 2 points (x1,y1) and (x2,y2) are given = \(\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (jdoe0001):

hmm rather \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 2}})\quad &({\color{red}{ 5}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (akashdeepdeb):

How are you soooo good at LaTeX? O.O

OpenStudy (anonymous):

so could you guys give me the answer for just this one

OpenStudy (anonymous):

1\7